# Weak interaction axial form factors of the octet baryons in nuclear medium

### v1

@@ V.4 Discussion
The results of the previous sections for $G_{A}$ and $G_{P}$, in vacuum and in the medium, are a natural consequence of the formalism proposed and discussed in Secs. III and IV. As a consequence, the properties of the functions $G_{A}$ and $G_{P}$ are the byproduct of the assumptions and methods used. We discuss here the virtues and the limitations of the present calculations.
@@
Some of the results for $G_{A}$ and $G_{P}$ mentioned in the previous sections are the outcome of the methodology used to determine the free parameters of the model to the lattice QCD data for the nucleon [8, 38]. The calibration based on the lattice QCD data has limited by the lack of data between $Q^{2}=0$ and $Q^{2}=0.15$ GeV2. This lack of data has consequences on the determination of the $n_{P}$ and on the quark form factor $g_{P}^{q}$, and may affect the accuracy of the calibration of the functions $G_{A}$ and $G_{P}$ at low $Q^{2}$.
@@
We first comment on the results for $G_{A}$. As mentioned already, the weak axial-vector form factor is very smooth near $Q^{2}=0$. This property was already present in the model for the vacuum [8], but was not so clearly observed due to the competition between the bare contribution and the meson cloud contribution with a sharp shape $\propto 1/(1+Q^{2}/M_{A}^{2})^{4}$. The smooth shape of the bare contribution can be seen in the bare contribution of Fig. 2 for the nucleon, and it is a consequence of $P$-wave contributions. Since the $P$-state mixture was determined by a fit to lattice and physical data in a range $Q^{2}=0$–3 GeV2 (equal weight for all $Q^{2}$ points), no special constraint was included in the fit to the data near $Q^{2}=0$. This is the main reason why the square axial radius deviates from the estimate based on a dipole function. In medium, the function $G_{A}$ has become softer when $|n_{P}|$ increases. The present results for $G_{A}$ are then mainly a consequence of the determination of the parameters of the model based on the lattice QCD simulations, where the low-$Q^{2}$ region is not well represented, due to the gap of 0.3, 0.4 GeV2 between the lattice QCD data points [38]. The properties of the function $G_{A}$ for the nucleon are extended to the octet baryon case.
@@
Also related to the previous topic are the results for $G_{P}$ at low $Q^{2}$. Contrarily to the functions $G_{A}$, there are no lattice QCD data for $Q^{2}=0$ (as there are no lattice QCD data for the Pauli form factor $F_{2}$). Since the function $g^{q}_{P}(Q^{2})$, defined in Eq. (12), contributes only to the bare contribution in the induced pseudoscalar form factor $G_{P}$ [see Eq. (18)], the free parameters of $g^{q}_{P}(Q^{2})$ are determined exclusively by the fit to the $G_{P}$ lattice QCD data. In these conditions, the calibration of the function $g_{P}^{q}$ is the result of the data above the minimum value for $Q^{2}$ ($Q^{2}>0.15$ GeV2) which may significantly affect the result for $G_{P}$ near $Q^{2}=0$, since that is the point where the global function $G_{P}$ is expected to have its maximum. More recent data with a wide range in $Q^{2}$, or independent estimates of $G_{P}(0)$ may help to improve the calibration of $G_{P}^{\rm B}$ for $\Delta I=1$ and $\Delta S=1$. In the alternative, a fitting procedure with some theoretical constraints to the function $G_{P}$ at low $Q^{2}$ may also be used in future studies.
@@
The parametrization of the meson cloud contributions at the level of the electromagnetic structure can also be improved with the inclusion of terms associated with the kaon cloud for the $\Lambda$, $\Sigma$ and $\Xi$ systems. The inclusion of the new terms will be sufficient to improve the description of the $\Sigma$ and $\Xi$ weak decay data, because it reduces the value of $\sqrt{Z_{B}}$, and consequently the magnitude of the $G_{A}$ (see Sec. V.1).
In the future, we may also consider a less phenomenological and more physical motivated description of the meson cloud contribution for the axial form factors, instead of the simplified SU(3) baryon-meson interaction (functions $F$ and $D$). That can be done considering a microscopic connection between the covariant spectator quark model and the bag model relating the electromagnetic-baryon and meson-baryon couplings of the two frameworks as in Refs. [61, 62].
@@ VI Numerical results for neutrino-nucleus cross sections
In the present section, we exemplify how the weak axial form factors calculated in the previous section, can be used to calculate neutrino-nucleus and antineutrino-nucleus cross sections considering the interaction of a neutrino or antineutrino with nucleons immersed in a nuclear matter with density $\rho$. In addition to the form factor $G_{A}$ we need also to know the electric ($G_{E}$) and magnetic ($G_{M}$) form factors in-medium. In that case we consider the model from Ref. [28] also derived within the covariant spectator quark model formalism.
The interactions with neutrinos/antineutrinos can be divided in neutral current (NC) and charged current (CC) reactions where leptons ($e,\mu$) can be found in the final state. Here we use $M$ to represent the nucleon mass in-vacuum or in-medium. For convenience, we analyze the reaction at the Lab frame, where the nucleon (momentum $p$) is at rest and the neutrino/antineutrino (momentum $k$) has an energy $E_{\nu}$, and the transfer momentum to the nucleon is $q$. In these conditions, the transition cross sections can be expressed in terms of the invariants $Q^{2}$, $y=\frac{p\cdot q}{p\cdot k}=\frac{Q^{2}}{2ME_{\nu}}$, the energy $E_{\nu}$, and the nucleon electromagnetic and axial form factors, as described below.
@@
where $G_{F}=1.664\times 10^{-5}$ GeV-2 is the Fermi constant, and $G_{\ell}^{\rm NC}$ ($\ell=E,M$), $G_{A}^{\rm NC}$ are the NC form factors, as defined next. The upper sign in the last term corresponds to the reactions with neutrinos and the lower sign the reaction with antineutrinos.
@@
Again, the signs in the last term respectively correspond to the reactions with neutrinos (upper sign) or antineutrinos (lower sign).
@@
where $M^{\prime}=M_{N}^{\ast}+V$ and $V$ is a vector potential term associated with the nucleon-medium interaction. Numerically one has $V=0.0627$ GeV for $\rho=0.5\rho_{0}$ and $V=0.1253$ GeV for $\rho=\rho_{0}$. The last correction is a consequence of the modification of the nucleon energy at rest in-medium, where the nucleon momentum take the form $p=(M^{\prime},{\bf 0})$ [23]. The transfer momentum $q$ is then modified in-medium [^1] The transfer momentum takes the form $q=(\omega,{\bf q})$, where $\omega=\frac{Q^{2}}{2M^{\prime}}$ and $|{\bf q}|^{2}=Q^{2}\left(1+\frac{Q^{2}}{4M^{\prime 2}}\right)$. leading to the result (49).
@@ VI.1 Form factors for NC and CC reactions
The form factors for NC transitions are defined by Eqs. (42) and (45) for proton and neutron targets, and are presented at the top (proton) and central (neutron) panels of Fig. 13, respectively. The form factors for the CC transitions are defined by Eqs. (47) and (48) and are presented at the bottom panel of Fig. 13.
The results from Fig. 13 exhibit a reduction of the magnitude in the form factors (absolute values) in the nuclear medium. The reduction increases with the density $\rho$ (the form factors are quenched in-medium). These results are a consequence of the combination of proton and neutron form factors. As discussed in Refs. [28, 29] proton and neutron form factors are modified differently in the nuclear medium. The exceptions to the suppression are the results of the electric and the magnetic form factors at $Q^{2}=0$. The suppression of $G_{A}$ in medium was already discussed in Sec. IV.
At $Q^{2}=0$ the electric form factor is unmodified by the medium because the electric charge does not change in-medium. As for the magnetic form factor at $Q^{2}=0$, the invariance of $G_{M}^{NC}(0)$ is only approximated, and it is a consequence of the general dominance of the valence quark contributions in the nuclear medium [28]. [The quark anomalous magnetic moments in natural units are not modified in medium.]
The medium effect on the magnetic form factors has a different trend when expressed in terms of the nuclear magneton in-vacuum [units $e/(2M_{N})$], instead of the natural units, as discussed in Ref. [28]. In the first case the nucleon magnetic form factors in-medium are corrected by the factor $\frac{M_{N}}{M_{N}^{\ast}}$ [28]. The consequence of this conversion is that since $G_{M}^{\rm NC}(0)$ in natural units is almost independent of the medium, $G_{M}^{\rm NC}(0)$ is enhanced when expressed in units $e/(2M_{N})$.
For the cases under discussion, the functions $G_{M}^{\rm NC}(Q^{2})$ and $G_{M}^{\rm CC}(Q^{2})$ from Fig. 13 (combinations of proton and neutron form factors), we can conclude that the amplification due to the factor $\frac{M_{N}}{M_{N}^{\ast}}$ dominate over the suppression of $G_{M}^{\rm NC}(Q^{2})$ in natural units. In conclusion, $G_{M}(Q^{2})$ is enhanced in-medium at low $Q^{2}$, when we use units $e/(2M_{N})$. When expressed in natural units, $G_{M}^{\rm NC}(Q^{2})$ and $G_{M}^{\rm NC}(Q^{2})$ are suppressed in-medium.
The results from Fig. 13 show that the function $G_{E}^{\rm NC}$ is very small for proton targets, as a consequence of the small magnitude of $1-4\sin^{2}\theta_{\rm w}$. For the discussion, we also notice that the magnitude of the form factors $G_{E}$ and $G_{A}$ for CC reactions are about two times the magnitudes for NC reactions with neutron targets. This effect has an impact on the magnitude of the neutrino-nucleon cross sections, as discussed later.
@@ VI.2 Kinematics
neglecting the neutrino mass. The same analysis is valid for antineutrino reactions. From the previous relation we can conclude that range of values of $Q^{2}$ according to the possible values of $u=\cos\theta_{l}$: $-1\leq u\leq 1$ are in the interval
@@ VI.3 Cross sections for NC reactions
The NC transitions for proton and neutron targets are presented in Figs. 14 and 15, respectively, for the densities $\rho=0$, $0.5\rho_{0}$ and $\rho_{0}$. In the upper panels we represent the $\nu N\to\nu N$ transitions and in the central panels the $\bar{\nu}N\to\bar{\nu}N$ transitions. The vertical lines represent the upper limit of $Q^{2}$ associated with the neutrino energies $E_{\nu}=5$, 20 and 100 MeV.
The first conclusion from the observations of the figures is that the single-differential cross sections are reduced in-medium (quenched effect) and that the suppression is more significant for higher densities. These results are a consequence of the suppression of the form factors $G_{\ell}^{\rm NC}$ and $G_{A}^{\rm NC}$ in-medium, discussed earlier. The difference of magnitudes between proton and neutron reactions is a consequence of the magnitude of the form factors [see Fig. 13].
@@
For a better visualization of this effect, we compare at the bottom panel of Figs. 14 and 15 the results for $\nu N\to\nu N$ the $\bar{\nu}N\to\bar{\nu}N$ for a fixed density ($\rho=0.5\rho_{0}$). Similar results are obtained for $\rho=0$ and $\rho=\rho_{0}$.
In the limit $Q^{2}=0$ the neutrino and antineutrino cross sections have the same value because the last terms of Eqs. (39) and (46) vanishes. The variation of the antineutrino cross sections with $Q^{2}$: almost constant for proton targets and the reduction with neutron targets is a consequence of the dependence of $G_{M}^{\rm NC}$ with $Q^{2}$. The product $G_{M}^{\rm NC}G_{A}^{\rm NC}$ is negative for protons and neutrons but the magnitude is larger for the case of the neutrons, because $|G_{M}^{\rm NC}|$ is larger for neutron targets ($|G_{A}^{\rm NC}|$ is the same for both cases). As a consequence the antineutrinos single-differential cross section is reduced for the neutron target when $Q^{2}$ increases.
@@ VI.4 Cross sections for CC reactions
The single-differential cross sections for the CC reactions $\nu n\to e^{-}p$ and $\bar{\nu}p\to e^{+}n$, are presented in Fig. 16 for the densities $\rho=0$, $0.5\rho_{0}$ and $\rho_{0}$. We consider the range of $Q^{2}$ associated to the neutrino energies $E_{\nu}=5$, 20 and 100 MeV.
As for the NC transitions, we conclude that the cross sections are reduced in-medium (quenched effect). Also for NC transitions one can observe the dominance of the neutrino cross sections over the antineutrino cross sections.
In comparison with the NC transitions we can notice that the scale of the single-differential cross sections increases by a factor of about four. This result is a consequence of the relations between the form factors $G_{E}$ and $G_{A}$ for NC reactions with neutron targets and CC reactions, discussed previously. The form factors for CC reactions are about two times larger. The effect of $G_{M}$ in the cross sections is reduced by the multiplication with $y$.

### v2

@@ V.5 Discussion
The results of the previous sections for $G_{A}$ and $G_{P}$, in vacuum and in the medium, are a natural consequence of the formalism proposed and discussed in Secs. III and IV. As a consequence, the properties of the functions $G_{A}$ and $G_{P}$ are the by-product of the assumptions and methods used. We discuss here the virtues and the limitations of the present calculations.
@@
Some of the results for $G_{A}$ and $G_{P}$ mentioned in the previous sections are the outcome of the methodology used to determine the free parameters of the model to the lattice QCD data for the nucleon [7, 38]. The calibration based on the lattice QCD data has been limited by the lack of data between $Q^{2}=0$ and $Q^{2}=0.15$ GeV2. This lack of data has consequences on the determination of the $n_{P}$ and on the quark form factor $g_{P}^{q}$, and may affect the accuracy of the calibration of the functions $G_{A}$ and $G_{P}$ at low $Q^{2}$.
@@
We first comment on the results for $G_{A}$. As mentioned already, the weak axial-vector form factor is very smooth near $Q^{2}=0$. This property was already present in the model for the vacuum [7], but was not so clearly observed due to the competition between the bare contribution and the meson cloud contribution with a sharp shape $\propto 1/(1+Q^{2}/M_{A}^{2})^{4}$. The smooth shape of the bare contribution can be seen in the bare contribution in Fig. 2 for the nucleon, and it is a consequence of $P$-wave contributions. Since the $P$-state mixture was determined by a fit to lattice and physical data in a range $Q^{2}=0$–3 GeV2 (equal weight for all $Q^{2}$ points), no special constraint was included in the fit to the data near $Q^{2}=0$. This is the main reason why the square axial radius deviates from the estimate based on a dipole function. In medium, the function $G_{A}$ has become softer when $|n_{P}|$ increases. The present results for $G_{A}$ are then mainly a consequence of the determination of the parameters of the model based on the lattice QCD simulations, where the low-$Q^{2}$ region is not well represented, due to the gap of 0.3 or 0.4 GeV2 between the lattice QCD data points [38]. The properties of the function $G_{A}$ for the nucleon are extended to the octet baryon case.
@@
Also related to the previous topic are the results for $G_{P}$ at low $Q^{2}$. Contrarily to the functions $G_{A}$, there are no lattice QCD data for $Q^{2}=0$ (as there are no lattice QCD data for the Pauli form factor $F_{2}$). Since the function $g^{q}_{P}(Q^{2})$, defined in Eq. (12), contributes only to the bare contribution in the induced pseudoscalar form factor $G_{P}$ [see Eq. (18)], the free parameters of $g^{q}_{P}(Q^{2})$ are determined exclusively by the fit to the $G_{P}$ lattice QCD data. In these conditions, the calibration of the function $g_{P}^{q}$ is the result of the data above the minimum value for $Q^{2}$ ($Q^{2}>0.15$ GeV2) which may significantly affect the result for $G_{P}$ near $Q^{2}=0$, since that is the point where the global function $G_{P}$ is expected to have its maximum. More recent data with a wide range in $Q^{2}$, or independent estimates of $G_{P}(0)$ may help to improve the calibration of $G_{P}^{\rm B}$ for $|\Delta I|=1$ and $|\Delta S|=1$. In the alternative, a fitting procedure with some theoretical constraints to the function $G_{P}$ at low $Q^{2}$ may also be used in future studies.
@@
The parametrization of the meson cloud contributions at the level of the electromagnetic structure can also be improved with the inclusion of terms associated with the kaon cloud for the $\Lambda$, $\Sigma$ and $\Xi$ systems. The inclusion of the new terms will be sufficient to improve the description of the $\Sigma$ and $\Xi$ weak decay data, because it reduces the value of $\sqrt{Z_{B}}$ and, consequently, the magnitude of the $G_{A}$ (see Sec. V.1).
In the future, we may also consider a less phenomenological and more physical motivated description of the meson cloud contribution for the axial form factors, instead of the simplified SU(3) baryon-meson interaction (functions $F^{\prime}$ and $D^{\prime}$). That can be done considering a microscopic connection between the covariant spectator quark model and the bag model relating the electromagnetic-baryon and meson-baryon couplings of the two frameworks as in Refs. [61, 62].
@@ VI Numerical results for neutrino-nucleus cross sections
In the present section, we exemplify how the weak axial form factors calculated in the previous section, can be used to calculate neutrino-nucleus and antineutrino-nucleus cross sections considering the interaction of a neutrino or antineutrino with nucleons immersed in a nuclear matter with density $\rho$. In addition to the form factor $G_{A}$ we need also to know the electric ($G_{E}$) and magnetic ($G_{M}$) form factors in medium. In that case we consider the model from Ref. [28] also derived within the covariant spectator quark model formalism.
The interactions with neutrinos/antineutrinos can be divided in neutral current (NC) and charged current (CC) reactions where leptons ($e,\mu$) can be found in the final state. Here we use $M$ to represent the nucleon mass in vacuum or in medium. For convenience, we analyze the reaction at the Lab frame, where the nucleon (momentum $p$) is at rest and the neutrino/antineutrino (momentum $k$) has an energy $E_{\nu}$, and the transfer momentum to the nucleon is $q$. In these conditions, the transition cross sections can be expressed in terms of the invariants $Q^{2}$, $y=\frac{p\cdot q}{p\cdot k}=\frac{Q^{2}}{2ME_{\nu}}$, the energy $E_{\nu}$, and the nucleon electromagnetic and axial form factors, as described below.
@@
where $G_{F}=1.664\times 10^{-5}$ GeV-2 is the Fermi constant, and $G_{\ell}^{\rm NC}$ ($\ell=E,M$) and $G_{A}^{\rm NC}$ are the NC form factors, as defined next. The upper sign in the last term corresponds to the reactions with neutrinos and the lower sign the reaction with antineutrinos.
@@
Again, the signs in the last term, respectively, correspond to the reactions with neutrinos (upper sign) or antineutrinos (lower sign).
@@
where $M^{\prime}=M_{N}^{\ast}+V$ and $V$ is a vector potential term associated with the nucleon-medium interaction. Numerically one has $V=0.0627$ GeV for $\rho=0.5\rho_{0}$ and $V=0.1253$ GeV for $\rho=\rho_{0}$. The last correction is a consequence of the modification of the nucleon energy at rest in medium, where the nucleon momentum takes the form $p=(M^{\prime},{\bf 0})$ [23]. The transfer momentum $q$ is then modified in medium [^1] The transfer momentum takes the form $q=(\omega,{\bf q})$, where $\omega=\frac{Q^{2}}{2M^{\prime}}$ and $|{\bf q}|^{2}=Q^{2}\left(1+\frac{Q^{2}}{4M^{\prime 2}}\right)$. leading to the result (49).
@@ VI.1 Form factors for NC and CC reactions
The form factors for NC transitions are defined by Eqs. (42) and (45) for proton and neutron targets, and are presented at the top (proton) and central (neutron) panels in Fig. 11, respectively. The form factors for the CC transitions are defined by Eqs. (47) and (48) and are presented at the bottom panel in Fig. 11.
The results from Fig. 11 exhibit a reduction of the magnitude in the form factors (absolute values) in the nuclear medium. The reduction increases with the density $\rho$ (the form factors are quenched in medium). These results are a consequence of the combination of proton and neutron form factors. As discussed in Refs. [28, 29] proton and neutron form factors are modified differently in the nuclear medium. The exceptions to the suppression are the results of the electric and the magnetic form factors at $Q^{2}=0$. The suppression of $G_{A}$ in medium was already discussed in Sec. IV.
At $Q^{2}=0$ the electric form factor is unmodified by the medium because the electric charge does not change in medium. As for the magnetic form factor at $Q^{2}=0$, the invariance of $G_{M}^{\rm NC}(0)$ is only approximated, and it is a consequence of the general dominance of the valence quark contributions in the nuclear medium [28]. [The quark anomalous magnetic moments in natural units are not modified in medium.]
The medium effect on the magnetic form factors has a different trend when expressed in terms of the nuclear magneton in vacuum [units $e/(2M_{N})$], instead of the natural units, as discussed in Ref. [28]. In the first case the nucleon magnetic form factors in medium are corrected by the factor $\frac{M_{N}}{M_{N}^{\ast}}$ [28]. The consequence of this conversion is that since $G_{M}^{\rm NC}(0)$ in natural units is almost independent of the medium, $G_{M}^{\rm NC}(0)$ is enhanced when expressed in units $e/(2M_{N})$.
For the cases under discussion, the functions $G_{M}^{\rm NC}(Q^{2})$ and $G_{M}^{\rm CC}(Q^{2})$ from Fig. 11 (combinations of proton and neutron form factors), we can conclude that the amplification due to the factor $\frac{M_{N}}{M_{N}^{\ast}}$ dominates over the suppression of $G_{M}^{\rm NC}(Q^{2})$ in natural units. In conclusion, $G_{M}(Q^{2})$ is enhanced in medium at low $Q^{2}$, when we use units $e/(2M_{N})$. When expressed in natural units, $G_{M}^{\rm NC}(Q^{2})$ and $G_{M}^{\rm NC}(Q^{2})$ are suppressed in medium.
The results from Fig. 11 show that the function $G_{E}^{\rm NC}$ is very small for proton targets, as a consequence of the small magnitude of $1-4\sin^{2}\theta_{\rm w}$. For the discussion, we also notice that the magnitude of the form factors $G_{E}$ and $G_{A}$ for CC reactions are about 2 times the magnitudes for NC reactions with neutron targets. This effect has an impact on the magnitude of the neutrino-nucleon cross sections, as discussed later.
@@ VI.2 Kinematics
neglecting the neutrino mass. The same analysis is valid for antineutrino reactions. From the previous relation we can conclude that the range of values of $Q^{2}$ according to the possible values of $u=\cos\theta_{l}$: $-1\leq u\leq 1$ is in the interval
@@ VI.3 Cross sections for NC reactions
The NC transitions for proton and neutron targets are presented in Figs. 12 and 13, respectively, for the densities $\rho=0$, $0.5\rho_{0}$ and $\rho_{0}$. In the upper panels we represent the $\nu N\to\nu N$ transitions and in the central panels the $\bar{\nu}N\to\bar{\nu}N$ transitions. The vertical lines represent the upper limit of $Q^{2}$ associated with the neutrino energies $E_{\nu}=5$, 20 and 100 MeV. The results for the different energies can be inferred from the upper limit for $Q^{2}$.
The first conclusion from the observations of the figures is that the single-differential cross sections are reduced in medium (quenched effect) and that the suppression is more significant for higher densities. These results are a consequence of the suppression of the form factors $G_{\ell}^{\rm NC}$ and $G_{A}^{\rm NC}$ in medium, discussed earlier. The difference of magnitudes between proton and neutron reactions is a consequence of the magnitude of the form factors [see Fig. 11].
@@
For a better visualization of this effect, we compare at the bottom panel in Figs. 12 and 13 the results for $\nu N\to\nu N$ the $\bar{\nu}N\to\bar{\nu}N$ for a fixed density ($\rho=0.5\rho_{0}$). Similar results are obtained for $\rho=0$ and $\rho=\rho_{0}$.
In the limit $Q^{2}=0$ the neutrino and antineutrino cross sections have the same value because the last terms of Eqs. (39) and (46) vanishes. The magnitudes of the cross sections are larger for $E_{\nu}=20$ MeV. The variation of the antineutrino cross sections with $Q^{2}$ for larger values of $E_{\nu}$: almost constant for proton targets and a soft reduction with neutron targets is a consequence of the dependence of $G_{M}^{\rm NC}$ with $Q^{2}$. The variation of the antineutrino cross sections with $Q^{2}$: almost constant for proton targets and the reduction with neutron targets is a consequence of the dependence of $G_{M}^{\rm NC}$ with $Q^{2}$. The product $G_{M}^{\rm NC}G_{A}^{\rm NC}$ is negative for protons and neutrons but the magnitude is larger for the case of the neutrons, because $|G_{M}^{\rm NC}|$ is larger for neutron targets ($|G_{A}^{\rm NC}|$ is the same for both cases). As a consequence, the antineutrino single-differential cross section is reduced for the neutron target when $Q^{2}$ increases.
@@ VI.4 Cross sections for CC reactions
The single-differential cross sections for the CC reactions $\nu n\to e^{-}p$ and $\bar{\nu}p\to e^{+}n$ are presented in Fig. 14 for the densities $\rho=0$, $0.5\rho_{0}$ and $\rho_{0}$. We consider the range of $Q^{2}$ associated to the neutrino energies $E_{\nu}=5$, 20 and 100 MeV. As before, the results for the different energies can be inferred from the upper limit lines for $Q^{2}$.
As for the NC transitions, we conclude that the cross sections are reduced in medium (quenched effect). Also for NC transitions one can observe the dominance of the neutrino cross sections over the antineutrino cross sections.
In comparison with the NC transitions we can notice that the scale of the single-differential cross sections increases by a factor of about 4. This result is a consequence of the relations between the form factors $G_{E}$ and $G_{A}$ for NC reactions with neutron targets and CC reactions, discussed previously. The form factors for CC reactions are about 2 times larger. The effect of $G_{M}$ in the cross sections is reduced by the multiplication with $y$.